OpenStudy (anonymous):

Which of the following points is on the graph of y = log1/2x?

OpenStudy (anonymous):

(1, 1/2) (1/2, 1) (0, 1)

OpenStudy (jdoe0001):

well, you know that a logarithmic function it's just another form of an exponential one that is \(\large y^\color{red}{n} = m\) in log notation you'd say \(\large log_ym = \color{red}{n}\)

OpenStudy (anonymous):

I don't understand at all

OpenStudy (jdoe0001):

well, ahemm, that'd be the notation, I'd think you would have covered this by now, but anyhow gimme a number raised to some power and its result I'll put it in logarithmic notation, just to show you

OpenStudy (jdoe0001):

say \(\large 2^\color{red}{3} = 8\) \( \large \implies log_28 = \color{red}{3}\)

OpenStudy (jdoe0001):

see the relation?

OpenStudy (anonymous):

I understand that, I just dont see how to find the answer above?

OpenStudy (jdoe0001):

when the subscript for the log is omitted, then is implicitly understood is base 10 so y = log1/2x really means \(\huge y =log_{10}\pmatrix{\frac{1}{2x}}\)

OpenStudy (anonymous):

so 1, 1/2

OpenStudy (jdoe0001):

so let's look at your numbers and test them about (1, 1/2) (1/2, 1) (0, 1) if x = 1, our fraction will turn to 1/2x = > 1/2(1) = 1/2 so what is the "exponent" we need to raise 10 to get 1/2? well, if we use the "1" given there, 10^1 = 10

OpenStudy (jdoe0001):

woops, I meant... \(10^{\frac{1}{2}}\)

OpenStudy (jdoe0001):

do you have a 4th choice?

OpenStudy (jdoe0001):

is it \(\huge y =log_{10}\pmatrix{\frac{1}{2}x} \)instead?

OpenStudy (anonymous):

here is different choices: x = 2y x = 2-y x = -(2y)

OpenStudy (jdoe0001):

hmmm, I don't see any of those choices as available